OpenStudy (anonymous):

x^3-x-1=0,find the value of

OpenStudy (anonymous):

\[\frac{ 1-a }{ 1+a }+\frac{ 1-b }{ 1+b}+\frac{ 1-c }{ 1+c }\] if \[a,b,c \] are roots of the equation

OpenStudy (anonymous):

\[\frac{ 1 }{ 1+a }+\frac{ 1 }{ 1+b }+\frac{ 1 }{ 1+c }-3\] if we sibstitute all roots individually\[a^3=a+1,b^3=b+1,c^3=c+1\] hence \[\frac{ 1}{ a^3 }+\frac{ 1 }{ b^3 }+\frac{ 1 }{ c^3 }-3\]